# The Effects of Multi-Task Learning on ReLU Neural Network Functions

### v2

@@ 4 Multivariate Multi-task Neural Network Training
for some dataset $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{N},{\bm{y}}_{N})\in\mathbb{R}^{d}\times\mathbb{R}^{T}$, where ${\mathcal{L}}$ is any convex loss function which is lower semicontinuous in its second argument and separable across the $T$ tasks [^4] In other words, ${\mathcal{L}}$ is expressible as ${\mathcal{L}}({\bm{u}},{\bm{v}})=\sum_{t=1}^{T}\tilde{{\mathcal{L}}}(u_{t},v_{t})$ for any ${\bm{u}},{\bm{v}}\in\mathbb{R}^{T}$, for some univariate loss function $\tilde{{\mathcal{L}}}$. For notational convenience, we denote both the multivariate and univariate loss functions as ${\mathcal{L}}$., and $K\geq N^{2}$.
We are interested in analyzing the behavior of solutions to (11) as the number of tasks $T$ grows. If $T$ is very large and the tasks are “diverse” (meaning informally that no single task has a strong influence on any other), one might intuitively expect that each neuron of an optimal solution to (11) would contribute to many of the tasks, and thus that output weight $v_{ks}^{*}$ for an individual neuron $k$ and task $s$ would be relatively small compared to the sum of the output weights $v_{kt}^{*}$ for tasks $t\neq s$. In this case, the $k^{\textrm{th}}$ term of the regularizer in (11) would be approximately equal to
@@
The above reasoning can be made precise. To formalize the notion of task diversity, we use the concept of exchangeability. A sequence $X_{1},\dots,X_{n}$ of random variables is exchangeable if their joint distribution remains unchanged for any permutation of the indices $1,\dots,n$. Intuitively, exchangeability captures the notion that the order of the random variables doesn’t matter; this is naturally consistent with most multi-task learning problems, in which the order in which the tasks are enumerated is irrelevant (it does not matter whether any particular task is labeled as task 1, task 2, etc.). Exchangeability can also be viewed as a weaker notion of independence of random variables: i.i.d. random variables are necessarily exchangeable, although the converse is not necessarily true. Another example of a random process generating exchangeable task labels is the following: let $\mu$ be a probability measure over a family of conditional probability distributions ${\cal P}$. Suppose that $p_{1},\dots,p_{T}\overset{iid}{\sim}\mu$ and ${\bm{y}}_{:t}|{\bm{x}}_{1},\dots,{\bm{x}}_{N}\sim p_{t}$, where ${\bm{y}}_{:t}:=[y_{1t},\dots,y_{Nt}]$ is the vector of labels corresponding to task $t$. Then ${\bm{y}}_{:1},\dots,{\bm{y}}_{:T}$ are exchangeable, conditioned on the data points ${\bm{x}}_{1},\dots,{\bm{x}}_{N}$, [^5] In the remainder of the discussion, all statements are interpreted as being conditioned on ${\bm{x}}_{1},\dots,{\bm{x}}_{N}$..
@@
over ${\mathbb{R}}^{K}$, where $s$ is an individual task. $J$ is simply the objective function of (11) with all parameters except for $v_{1s},\dots,v_{Ks}$ held fixed at their optimal values, and with the (constant) data fitting terms from the other tasks except $s$ removed. Loosely speaking, we can view $J$ as the objective function of (11) from the perspective of a single task $s$. Note that the optimal values $v_{1s}^{*},\dots,v_{Ks}^{*}$ for (11) also minimize $J$; otherwise they would not be optimal for (11). The following theorem shows that, when the task labels are exchangeable, $J$ is approximately quadratic near its minimizer and that this approximation get stronger as the number of tasks $T$ increases.
@@
Suppose that ${\bm{y}}_{:1},\dots,{\bm{y}}_{:T}$ are exchangeable and let $\mathcal{S}$ be the set of optimal active neurons solving (11), that is, the neurons for which $\|{\bm{v}}^{*}_{k}\|_{2}>0$. For an individual task $s$, consider the objective
@@
The proof of Theorem 4.1 is in Appendix 6.2. The theorem states that the solution to the weighted $\ell^{2}$ minimization problem
@@
$\min_{v_{1},\dots,v_{K}}\sum_{i=1}^{N}{\mathcal{L}}\left(y_{i},\sum_{k=1}^{K}v_{k}\mathbf{\Psi}_{ik}\right)+\lambda\sum_{k=1}^{K}|v_{k}|.$ (18)
yields output weights which are exactly optimal for (11). Note that the matrices ${\bm{\Phi}}$ in 17 and $\mathbf{\Psi}$ in 18 are not the same, since they are determined by the optimal input weights and biases for (11), which are themselves data- and task-dependent. Nonetheless, comparing (17) and (18) highlights the different nature of solutions learned for (11) in the single-task versus multi-task case. The multi-task learning problem with exchangeable tasks favors linear combinations of the optimal neurons which have a minimal weighted $\ell^{2}$ regularization penalty. In contrast, the single-task learning problem favors linear combinations of optimal neurons which have a minimal $\ell^{1}$ penalty. Therefore, multi-task learning with a large number of diverse tasks promotes a fundamentally different linear combination of the optimal features learned in the hidden layer.
Additionally, exchangeability of the tasks allows for the following characterization of the coefficients $\gamma_{ks}$ in (17):
@@
The proof is in Section 6.2. This result implies that that the weights in the weighted $\ell^{2}$ regularizer for each task converge to a common value across all the tasks with high probability as $T$ grows larger. In particular, for large $T$
@@
$\langle f_{\bm{v}},f_{\bm{u}}\rangle_{\mathcal{H}}={\bm{v}}^{T}{\bm{Q}}{\bm{u}}$ (22)
@@
$\langle\kappa(\cdot,{\bm{x}}),f\rangle_{\mathcal{H}}=\left\langle\sum_{k=1}^{K}Q^{-1}_{kk}\phi_{k}({\bm{x}})\phi_{k},\sum_{k=1}^{K}v_{k}\phi_{k}\right\rangle_{\mathcal{H}}$ (24)
@@
$\langle\kappa(\cdot,{\bm{x}}),f\rangle_{\mathcal{H}}=\left\langle g_{{\bm{u}}},f\right\rangle_{\mathcal{H}}=\phi({\bm{x}}){\bm{Q}}^{-1}{\bm{Q}}{\bm{v}}=\sum_{k=1}^{K}v_{k}\phi({\bm{x}})=f({\bm{x}})$ (25)
@@
We provide empirical evidence for the claims presented in this section in Fig. 5 on a simple multi-variate dataset. First, we demonstrate the variety of solutions that interpolate this dataset in a single task setting. In contrast, we show that the solutions obtained via multi-task learning with additional random tasks are very similar and often much smoother than those obtained by single-task learning supporting our claim that these solutions are well approximated by a kernel method. We also verify that the optimization (14) is a good approximation for (13). We include additional experiments in Section 7 that demonstrate that these observations hold across multiple trials.
@@
Figure 5: Top Row: Three solutions to ReLU neural network interpolation (blue surface) of training data (red). The eight data points are located at the vertices of two squares, both centered at the origin. The outer square has side-length two and values of zero at the vertices. The inner square has side-length $1$ and values of one at the vertices. All three functions interpolate the data and are global minimizers of (2) and (3) when solving for just this task (i.e., $T=1$). Bottom Row: The left figure shows the solution to the first output of a multi-task neural network with $T=101$ tasks, the first is the original task depicted in the first row while the other $100$ tasks are randomly generated from a standard Gaussian distribution. The learned solutions are almost identical across multiple optimization runs, indicating that the solutions to the multi-task optimization are nearly unique. Here we show one representative example; more examples are depicted in Appendix 7 showing that this phenomenon holds across many runs. The figure on the right shows the solution to fitting the training data by solving (14) over a fixed set of features learned by a multi-task neural network trained to fit only the $T=100$ random tasks. We observe that unlike the highly variable solutions of single-task optimization problem, the solutions obtained by solving the multi-task optimizations are nearly identical, as one would have for kernel methods. Moreover, the solution obtained by solving (14) is also similar to the solution of the full multi-task training problem with all $T=101$ tasks. Due to the simplicity of this dataset the optimality of these solutions were confirmed by solving the equivalent convex optimization to (2) developed in Ergen and Pilanci (2021).
5 Conclusion and Discussion of Limitations/Future Work
@@
We have shown that univariate, multi-task shallow ReLU neural networks which are trained to interpolate a dataset with minimal sum of squared weights almost always represent a unique function. This function performs straight-line interpolation between consecutive data points for each task. This connect-the-dots solution is also the solution to a min-norm data-fitting problem in an RKHS. We provide mathematical analysis and numerical evidence suggesting that a similar conclusion may hold in the mulvariate-input case, as long as the tasks are sufficiently large in number and “diverse.” These results indicate that multi-task training of neural networks can produce solutions that are strikingly different from those obtained by single-task training, and highlights a novel connection between these multi-task solutions and kernel methods.

### v3

@@ 4 Multivariate Multi-Task Neural Network Training
for some dataset $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{N},{\bm{y}}_{N})\in\mathbb{R}^{d}\times\mathbb{R}^{T}$, where ${\mathcal{L}}$ is any loss function which is lower semicontinuous in its second argument and separable across the $T$ tasks [^5] In other words, ${\mathcal{L}}$ is expressible as ${\mathcal{L}}({\bm{u}},{\bm{v}})=\sum_{t=1}^{T}\tilde{{\mathcal{L}}}(u_{t},v_{t})$ for any ${\bm{u}},{\bm{v}}\in\mathbb{R}^{T}$, for some univariate loss function $\tilde{{\mathcal{L}}}$. For notational convenience, we denote both the multivariate and univariate loss functions as ${\mathcal{L}}$., and $K\geq N^{2}$ (see footnote 1).
We are interested in analyzing the behavior of solutions to (11) as the number of tasks $T$ grows. Intuitively, if $T$ is very large, it it is reasonable to expect that the optimal output weight $v_{ks}^{*}$ for an individual neuron $k$ and task $s$ would be relatively small compared to the sum of the output weights $v_{kt}^{*}$ for tasks $t\neq s$. In this case, the $k^{\textrm{th}}$ term of the regularizer in (11) would be approximately equal to
@@
The above reasoning can be made precise using the concept of exchangeability. A sequence of random variables is exchangeable if their joint distribution remains unchanged for any permutation of the sequence. Intuitively, exchangeability captures the notion that the order of the random variables does not matter. This is naturally consistent with multi-task learning problems, in which the order in which the tasks are enumerated and assigned to the outputs of the network is irrelevant. Therefore, the $T$ tasks can be assigned uniformly at random to the $T$ outputs of the network. This random assignment process induces a distribution on the training data (labels) for each output, denoted by ${\bm{y}}_{\cdot,1},\dots,{\bm{y}}_{\cdot,T}\in\mathbb{R}^{N}$. These vectors are comprised of the $N$ labels for each task corresponding to the inputs ${\bm{x}}_{1},\dots,{\bm{x}}_{N}$. In particular, ${\bm{y}}_{\cdot,1},\dots,{\bm{y}}_{\cdot,T}$ are exchangeable. In the remainder of the discussion, we assume that the vectors ${\bm{y}}_{\cdot,1},\dots,{\bm{y}}_{\cdot,T}$ have been generated according to this procedure.
@@
over ${\mathbb{R}}^{K}$, where $s$ is an individual task. $J$ is simply the objective function of (11) with all parameters except for $v_{1s},\dots,v_{Ks}$ held fixed at their optimal values, and with the (constant) data fitting terms from the other tasks except $s$ removed. Loosely speaking, we can view $J$ as the objective function of (11) from the perspective of a single task $s$. Note that the optimal values $v_{1s}^{*},\dots,v_{Ks}^{*}$ for (11) also minimize $J$; otherwise they would not be optimal for (11). The following theorem shows that, when the task labels are exchangeable, the regularization term in $J$ is approximately quadratic near its minimizer and that this approximation get stronger as the number of tasks $T$ increases.
@@
Let $\mathcal{S}$ be the set of optimal active neurons solving (11), that is, the neurons for which $\|{\bm{v}}^{*}_{k}\|_{2}>0$. For an individual task $s$, consider the objective
@@
The proof of Theorem 4.1 is in Section A.2. The theorem states that the solution to the followed weighted $\ell^{2}$ regularized problem
@@
$\min_{v_{1},\dots,v_{K}}\sum_{i=1}^{N}{\mathcal{L}}\left(y_{i},\sum_{k=1}^{K}v_{k}\mathbf{\Psi}_{ik}\right)+\lambda\sum_{k=1}^{K}|v_{k}|$ (18)
yields output weights which are exactly optimal for (11). Note that the matrices ${\bm{\Phi}}$ in (17) and $\mathbf{\Psi}$ in (18) are not the same, since they are determined by the optimal input weights and biases for (11), which are themselves data- and task-dependent. Nonetheless, comparing (17) and (18) highlights the different nature of solutions learned for (11) in the single-task versus multi-task case. The multi-task learning problem with exchangeable tasks favors linear combinations of the optimal neurons which have a minimal weighted $\ell^{2}$ regularization penalty. In contrast, the single-task learning problem favors linear combinations of optimal neurons which have a minimal $\ell^{1}$ penalty. Therefore, multi-task learning with a large number of tasks promotes a fundamentally different linear combination of the optimal features learned in the hidden layer.
Additionally, exchangeability of the tasks allows for the following characterization of the coefficients $\gamma_{ks}$ in (17) (cf Lemma A.3 in Section A.2):
@@
The proof is in Section A.2. This result suggests that the weights in the weighted $\ell^{2}$ regularizer for each task converge to a common value across all the tasks with high probability as $T$ grows larger. In particular, for large $T$:
@@
$\langle f_{\bm{v}},f_{\bm{u}}\rangle_{\mathcal{H}}={\bm{v}}^{\top}{\bm{Q}}{\bm{u}}$ (22)
@@
$\langle\kappa(\cdot,{\bm{x}}),f\rangle_{\mathcal{H}}=\left\langle\sum_{k=1}^{K}Q^{-1}_{kk}\phi_{k}({\bm{x}})\phi_{k},\sum_{k=1}^{K}v_{k}\phi_{k}\right\rangle_{\mathcal{H}}.$ (24)
@@
$\langle\kappa(\cdot,{\bm{x}}),f\rangle_{\mathcal{H}}=\left\langle g_{{\bm{u}}},f\right\rangle_{\mathcal{H}}=\phi({\bm{x}}){\bm{Q}}^{-1}{\bm{Q}}{\bm{v}}=\sum_{k=1}^{K}v_{k}\phi({\bm{x}})=f({\bm{x}}).$ (25)
@@
We provide empirical evidence for the claims presented in this section in Figure 5 on a simple multi-variate dataset. First, we demonstrate the variety of solutions that interpolate this dataset in a single task setting. In contrast, we show that the solutions obtained via multi-task learning with additional random tasks are very similar and often much smoother than those obtained by single-task learning supporting our claim that these solutions are well approximated by a kernel method. We also verify that the optimization (14) is a good approximation for (13). We include additional experiments in Appendix B that demonstrate that these observations hold across multiple trials.
(a) Solution to (11) with $T=1$ (single-task training)
(b) Solution to (11) with $T=1$ (single-task training)
(c) Solution to (11) with $T=1$ (single-task training)
(d) Solution to (11) with $T=101$
(multi-task training)
(e) Solution to (26)
(RKHS approximation)
@@
Figure 5: ReLU network interpolation in two-dimensions. The solutions shown were obtained with regularization parameter $\lambda\approx 0$. Top Row – Solutions to single-task training: Figures 5(a), LABEL:, 5(b), LABEL: and 5(c) show solutions to ReLU neural network interpolation (blue surface) of training data (red). The eight data points are located at the vertices of two squares, both centered at the origin. The outer square has side-length two and values of 00 at the vertices. The inner square has side-length one and values of $1$ at the vertices. All three functions interpolate the data and are global minimizers of (2) and (3) when solving for just this task (i.e., $T=1$). Due to the simplicity of this dataset the optimality of the solutions in the first row were confirmed by solving the equivalent convex optimization to (2) developed in Ergen and Pilanci (2021). Bottom Row – Solutions to multi-task training: Figure 5(d) shows the solution to the first output of a multi-task neural network with $T=101$ tasks. The first output is the original task depicted in the first row while the labels for other $100$ tasks are randomly generated i.i.d from a Bernoulli distribution with equal probability for one and zero. Here we show one representative example; more examples are depicted in Appendix B showing that this phenomenon holds across many runs. Figure 5(e) shows the solution to fitting the training data by solving (26) over a fixed set of features learned by the multi-task neural network with $T=100$ random tasks. We observe that unlike the highly variable solutions of single-task optimization problem, the solutions obtained by solving the multi-task optimizations are nearly identical, as one would have for kernel methods. Moreover, the solution obtained by solving (26) is also similar to the solution of the full multi-task training problem with all $T=101$ tasks.
5 Conclusion and Discussion
@@
We have shown that univariate, multi-task shallow ReLU neural networks which are trained to interpolate a dataset with minimal sum of squared weights almost always represent a unique function. This function performs straight-line interpolation between consecutive data points for each task. This solution is also the solution to a min-norm data-fitting problem in an RKHS. We provide mathematical analysis and numerical evidence suggesting that a similar conclusion may hold in the mulvariate-input case, as long as the tasks are sufficiently large in number. These results indicate that multi-task training of neural networks can produce solutions that are strikingly different from those obtained by single-task training, and highlights a novel connection between these multi-task solutions and kernel methods.